OpenStudy (anonymous):

square root of 2 + 4sinx cosx

OpenStudy (anonymous):

\[\sqrt{2}+4sinxcosx\]

OpenStudy (anonymous):

Try rewriting the former expression as \[2\frac{ \sqrt{2} }{2}+2\left( 2 \sin(x) \cos(x) \right)\], i.e. \[2\left(\frac{ \sqrt{2} }{2}+2 \sin(x) \cos(x) \right)\] ... then you can use \[\sin\left(\frac{\pi}{4}\right)=\frac{\sqrt{2}}{2}\] and \[2\sin(x)\cos(x)=\sin\left(2x\right)\] ... but the assignment itself is a bit unclear, maybe it was supposed to be an equation? Then it would make more sense.